# Video-Based Detection and Analysis of Errors in Robotic Surgical Training

### v1

@@ Abstract
Robot-assisted minimally invasive surgeries offer many advantages but require complex motor tasks that take surgeons years to master. There is currently a lack of knowledge on how surgeons acquire these robotic surgical skills. To help bridge this gap, we previously followed surgical residents learning complex surgical training dry-lab tasks on a surgical robot over six months. Errors are an important measure for self-training and for skill evaluation, but unlike in virtual simulations, in dry-lab training, errors are difficult to monitor automatically. Here, we analyzed the errors in the ring tower transfer task, in which surgical residents moved a ring along a curved wire as quickly and accurately as possible. We developed an image-processing algorithm to detect collision errors and achieved detection accuracy of $\sim 95\%$. Using the detected errors and task completion time, we found that the surgical residents decreased their completion time and number of errors over the six months. This analysis provides a framework for detecting collision errors in similar surgical training tasks and sheds light on the learning process of the surgical residents.
@@ I INTRODUCTION
In robot-assisted minimally invasive surgery (RAMIS) the surgeon uses surgeon-side manipulators to control a camera arm and patient-side manipulators, on which surgical instruments are mounted. RAMIS has become increasingly widespread over the past decades [1, 2], using systems such as the da Vinci surgical system (Intuitive Surgical Inc., Sunnyvale, California). RAMIS offers many benefits to both patients and surgeons, including less postoperative pain and tissue damage, shorter hospital stay [1, 3], reduced tremors compared to open surgeries [2], and improved ergonomics compared to laparoscopic surgeries [4, 2].
@@
To obtain these advantages, robotic surgeons require substantial training [5]. Although robotic surgical skill learning has been studied [6, 7, 8, 9, 10, 11, 12, 13, 14, 15], a long-term study on how surgeons acquire surgical skill was needed. Therefore, in [16], surgical residents came three times throughout a shift once a month for six months. They performed several surgical training tasks; video and kinematic data were recorded to study their learning processes. In the current work, we evaluate the performance in the ring tower transfer task, in which residents moved a ring along a curved wire as quickly as possible, while minimizing collisions with the wire.
@@
One challenge when assessing surgical skills is the method of evaluation. Currently, much of the evaluation is subjective, where one surgeon evaluates the performance of another and awards a score [17, 18, 19]. These scores suffer from large variability and require a great deal of time from the evaluating surgeons. RAMIS offers the possibility of overcoming these limitations by providing recorded kinematic and video data. This enables automatically computing metrics that are indicative of surgical skill [20, 21, 22, 23, 24, 12, 25, 26]. Metrics that have been used include task completion time [20, 22, 24], path length [21], and movement smoothness [22].
@@
Additionally, accuracy, or error metrics, can be important [22, 24], and are often assessed in motor control studies [27, 28, 29, 30]. Detecting errors is important as they have been shown to facilitate different learning processes [31], distinguishing between implicit learning driven by error in the predicted outcome of movement [32, 33, 34] and explicit learning driven by performance error [35, 34]. However, many of these motor control studies focus on simple tasks, and similar error metrics are challenging to compute in the more complex RAMIS tasks [13], leading many works to focus on other metrics, examples of which are listed above. Developing algorithms for detecting errors is important for eventually pinpointing the different learning mechanisms that underlie skill acquisition in robotic surgical training. In virtual simulations, analysis of errors is available, such as by Intuitive Surgical’s SimNow simulator [36]. However, detecting errors in dry-lab training is still an open challenge.
Several works have developed algorithms for error detection in surgical tasks [37, 38, 39, 40], which can enable computing error or accuracy metrics. For example, [37, 38, 39] trained neural networks to automatically detect errors in robotic surgical video data, such as the JIGSAWS dataset [41]. While these works demonstrate promising results, they require labeled datasets to train the networks.
@@
In this work, we aimed to develop an image-processing algorithm for detecting collision errors in the ring tower transfer task. This algorithm offers the advantage of not requiring labeled data and can be used to label the data for future analyses. We used the video data from [16], and wrote an algorithm to detect the collisions between the ring and the tower. Quantifying the errors can allow for studying their progression over the six-month experiment. As there is often a trade-off between speed and accuracy [24, 27, 29, 42], we assessed completion time together with the errors to shed light on the surgical residents’ learning process of the task. The main contributions of our work are:
@@
    We developed an image-processing algorithm for detecting collision errors between the ring and the tower.
@@
    We have released our codes and results, such that they can be adapted for similar tasks. Furthermore, all the labeled data will be released for future analyses.
@@
    We discuss the learning process of the residents over six months based on completion time and error results.
@@ II-A Dataset
18 surgical residents with no robotic experience from the Soroka Medical Center, Israel visited the lab 18 times – before, during, and after a 26-hour shift approximately once a month for six months. In each visit, they used the da Vinci Research Kit (dVRK) [43] to complete three surgical training tasks. They sat at the surgeon console and used the surgeon-side manipulators (SSMs) to teleoperate the patient-side manipulators (PSMs). The setup included two static Blackfly S cameras (FLIR Integrated Imaging Solutions Inc.) that captured the task board on the patient side. The video feed from the cameras was displayed on the stereo viewer in the surgeon console. The kinematic data, and video from the left camera, were recorded. In this work, we use the video data, which was presented to the residents at 35 Hz; however, the images were not recorded at a constant rate.
Here, we focus on the ring tower transfer task (Fig. 1); the residents extracted the black ring from the right vertical (RV) tower using their right hand, transferred it to their left hand, and inserted it onto the left horizontal (LH) tower. Next, they extracted the ring from the left vertical (LV) tower with their left hand and then inserted it onto the right horizontal (RH) tower using their right hand. The residents were asked to do so as fast as possible with minimal errors. Errors were defined as collisions between the tower and the ring. They received no feedback about their performance. Full details about the experiment and setup can be found in [16].
@@ II-B3 Error Detection
Each collision error of the ring with one of the towers was reflected as a movement of the tower in the recorded video. Therefore, to detect each error, our main steps were to: (1) locate the towers in the video frames, (2) compute all movement between each two consecutive frames, (3) sum the magnitude of the velocities of all pixels located as tower pixels, and (4) use a threshold to classify the error frames. We will now elaborate on each step of the algorithm.
@@
Computation of all movement between each two consecutive frames: We used the MATLAB optical flow function with the Horn-Schunck method [44]. From this, we obtained the magnitude of each of the image pixel velocities between every two consecutive images. This signal was likely affected by the lack of a constant sampling rate; however, it was still suitable for our needs.
@@
Lastly, we set several rules to clean the thresholded signal. The first was that a detected movement in the first 10 frames was only considered a movement if it continued also in the following five frames. This was because the first frame in every video was compared to a black image. Therefore, there was commonly a detected tower movement at the beginning, which was real in some cases, and an artifact in others. Next, all detected moves within 10 samples of each other were combined. This was to cope with cases in which the tower did not significantly move between two consecutive frames during an error. Third, all lone detected movement samples in a neighborhood of five samples were removed. Fourth, if there was a detected tower movement in the last 10 samples of an interaction with one of the vertical towers, the error lasted until the end of the interaction. We did not implement this for the horizontal towers, as the final part of the interaction was to place the ring on the tower, consistently leading to an intentional tower movement. Lastly, as part of the manual event segmentation, all cases of robot crashes were noted (this occurred a total of nine times during the interaction with one of the four towers), and the frames between the beginning of the crash and the return to the task were discarded from the data in all our analyses.
@@
Removal of End of Horizontal Tower Data: The residents placed the rings near the end of the horizontal towers. Hence, this was the only portion of the task in which a tower movement did not reflect an error, however, was still detected as one. Including this in the data introduces noise, and we therefore removed it as described in the following.
In an additional algorithm, we detected the black ring using a color threshold and a size threshold to prevent the detection of the much larger black robotic arm. Furthermore, as the ring is on the tower during the entire interaction with the tower, a minimal distance between each detected ring pixel and the nearest detected tower pixel was set. We do not elaborate on the details of this algorithm here, however, the codes and results are available in the github repository.
To identify when the residents moved to the straight part of the horizontal tower, where the ring was placed, we set a limit on the horizontal axis. This was because collisions with the horizontal towers led to tower movements primarily in the vertical axis. We set several such limits, creating separate tower segments, the last of which was the straight part at the end where residents were asked to place the ring. To avoid introducing noise into the data, we removed the data from this segment from the analyses.
Algorithm Analysis: Our algorithm outputted the frames with detected collisions for each resident’s interaction with each of the four towers. To check our algorithm, we displayed the results on video. Our code used the automatically detected tower movements and created a copy of the recorded task video while coloring the tower red every time a tower movement was detected (Video 1). We found that our algorithm worked well for most of the collisions. To account for occasions of missed, or wrongly detected collision errors, we wrote a code enabling us to run through the video and correct the automatic labels with a simple button press when necessary. This manual check enabled us to quantify our algorithm’s ability to accurately detect the collision errors.
To quantify the error detection ability of our algorithm, we compared the frames detected as error frames by the algorithm with any corrections made. We computed the number of True Positive (TP), True Negative (TN), False Positive (FP), and False Negative (FN) frames for each tower and each resident. Next, we computed the accuracy, true positive rate (TPR), true negative rate (TNR) and F1 score:
$Accuracy=\frac{TN+TP}{TN+FP+FN+TP}$ (1)
$TPR=\frac{TP}{TP+FN}$ (2)
$TNR=\frac{TN}{TN+FP}$ (3)
$F1=\frac{2TP}{2TP+FP+FN}$ (4)
@@ II-B4 Surgical Learning Analysis
where $T$ is the time it took to complete the insertion or extraction from each tower in [sec]. Each of these times was computed using the timestamps of the frames manually segmented as the start and end of the interaction with each tower. The timestamp of the start frame was subtracted from that of the end frame, and the times of the four towers were summed. Hence, the $Completion\;Time$ reflects the interaction time with the four towers, however, unlike in [16] does not include the time spent transferring the rings between hands or the positioning at the start of each tower, as these data were not included in the current work.
@@
where $ET$ is the total time spent colliding with each tower. The error time for each tower was computed by subtracting the first timestamp of the error from the last timestamp of the error. These times were then summed for each tower, and then for all four towers. This total error time was then divided by the $Completion\;Time$ to obtain the $Error\;Percentage$.
@@ II-B5 Statistical Analyses
We used three two-way repeated measures ANOVA models to assess the progression of each of our three metrics between months and sessions. In each model, the independent variables were shift (categorical: 1-6), timing relative to the shift (categorical: before, during, after), and resident (random). We also tested the interaction between the shift and the timing relative to the shift to assess if the effect of the timing differed between the shifts. The dependent variables were each the $Completion\;Time$, $Number\;of\;Errors$, and $Error\;Percentage$ in turn. For the $Completion\;Time$ and $Number\;of\;Errors$, we log-transformed the data to perform the analysis on normally distributed data. We used Mauchly’s test to check the assumption of sphericity and used Greenhouse-Geisser to adjust the degrees of freedom in the event of violations. The two participants with missing data were removed from the statistical analyses. We performed the statistical analyses using the MATLAB Statistics Toolbox. Statistical significance was determined at the 0.05 threshold.
@@ III Results
III-A Error Detection
Comparing the frames detected as errors by our algorithm with any manual corrections made allowed us to quantify the detection ability of our algorithm. We achieved an accuracy of 94.79%, a TPR (true positive rate, the ability to detect an error when there is one) of 91.98%, a TNR (true negative rate, the ability to recognize that there is no error when there indeed is not a collision) of 98.93%, and an F1 score of 95.46%. These values demonstrate the high ability of the algorithm to accurately detect the collision errors.
III-B Surgical Learning Analysis
The completion time and error results of all 18 residents are shown in Fig. 3. Fig. 3(a) shows the individual results of all the residents, as well as the average trend across the 18 visits. We observed a significant decrease in $Completion\;Time$, both between shifts and the timing relative to the shifts. This is further shown in Fig. 3(b) and Fig. 3(c), which are averaged across the shift, and the timing relative to the shift, respectively. This decrease reflects that, on average, the surgical residents learned to complete the task faster as the experiment progressed. The interaction between the shift and timing relative to the shift was also significant, indicating that the improvement rates in the $Completion\;Time$ between the three timings relative to the shift differed between the six shifts.
When examining the error metrics, we observed a relatively constant average $Error\;Percentage$. There was no significant difference between shifts, reflecting that residents spent a similar percentage of time colliding with the towers in the six months. This demonstrates a lack of improvement in terms of reducing errors. There was a significant effect of the timing relative to the shift on the $Error\;Percentage$, however, the effects were small (before: $65.12\pm 21.34\%$, during: $66.22\pm 22.56\%$, and after: $61.30\pm 23.03\%$). Lastly, the $Number\;of\;Errors$ significantly decreased between the six shifts, but not between the three timings relative to the shift. Taken together, the results show that the residents learned to complete the task faster, and began with making a larger number of shorter errors. With time, they made fewer separate errors, however, the portion of the time spent making the errors remained relatively constant. The statistical results are reported in Table III-B.
@@ III-B Surgical Learning Analysis
Figure 3: Time and Error Results. (a) The results of all 18 residents in the 18 visits are shown in pale purple. The dark line shows the average, and the shaded area is the 95% confidence interval, of all the residents. (b) The average results of all the residents as a function of the shift number. The three shades of green show the three timings relative to the shift (before, during and after). (c) The average results of all the residents as a function of the timing relative to the shift. The six shades of blue show the six shifts. The top row shows the $Completion\;Time$, the middle row is the $Error\;Percentage$ and the bottom row is the $Number\;of\;Errors$.
@@
Fig. 4 further illustrates the process in the time-error space. In terms of error metrics, we used the $Error\;Percentage$, as it describes the amount of task time spent in error, whereas the $Number\;of\;Errors$ does not reflect their magnitude in the task. Fig. 4(a) shows the average $Error\;Percentage$ of all the residents as a function of the average $Completion\;Time$ for each of the 18 visits. This illustrates the average improvement in the $Completion\;Time$ with a lack of deterioration in the $Error\;Percentage$.
@@
Fig. 4(b) shows each resident’s first and last $Error\;Percentage$ as a function of $Completion\;Time$, illustrating the individual process of each resident. We used a polynomial line to split the time-error space to demonstrate that all residents began on the right side of the line, and ended on the left. This line is similar to the Speed-Accuracy Tradeoff Functions shown in [29], characterizing the relation between the two. All the green symbols, representing the first visit, show a higher $Completion\;Time$ relative to the blue symbols (last visit), illustrating that all of the residents learned to complete the task faster over the six months. Many of the residents also showed some improvement in the $Error\;Percentage$, which taken together with the improvement in $Completion\;Time$ shows overall learning and improvement in both the time and error aspects of the task. However, there were six residents who improved their $Completion\;Time$ but increased their $Error\;Percentage$.
@@
Figure 4: Time-Error Analysis. (a) The average $Error\;Percentage$ as a function of the average $Completion\;Time$ across all 18 residents. The green shades indicate earlier sessions, and blue shades are the later ones, as shown in the color bar. (b) The $Error\;Percentage$ as a function of $Completion\;Time$ for each of the 18 residents in the first (green) and last (blue) visit. Each resident is shown by a different marker, and the first and last visits of each resident are connected with a line. To further demonstrate the improvement, the gray polynomial line splits the time-error space, such that all residents began on the right of the line and ended on the left.

### v2

# Analysis of Errors in Robotic Surgical Skill Acquisition with Video-Based Detection

@@ Abstract
Robot-assisted minimally invasive surgeries offer many advantages but require complex motor tasks that take surgeons years to master. There is currently a lack of knowledge on how surgeons acquire these robotic surgical skills. Toward bridging this gap, a previous study followed surgical residents learning complex surgical dry lab tasks on a surgical robot over six months. Errors are an important measure for training and skill evaluation, but unlike in virtual simulations, in dry lab training, errors are difficult to monitor automatically. Here, we analyzed errors in the ring tower transfer task, in which surgical residents moved a ring along a curved wire as quickly and accurately as possible. We developed an image-processing algorithm using color and size thresholds, optical flow and short time Fourier transforms to detect collision errors and achieved a detection accuracy of approximately 95%. Using the detected errors and task completion time, we found that the residents reduced their completion time and number of errors over the six months, while the percentage of task time spent making errors remained relatively constant on average. This analysis sheds light on the learning process of the residents and can serve as a step towards providing error-related feedback to robotic surgeons.
@@ I Introduction
In robot-assisted minimally invasive surgery (RAMIS), the surgeon uses surgeon-side manipulators to control a camera arm and patient-side manipulators, on which surgical instruments are mounted. RAMIS has become increasingly widespread over the past decades [hussain_use_2014, lanfranco_robotic_2004], using systems such as the da Vinci surgical system (Intuitive Surgical Inc., Sunnyvale, California). RAMIS offers many benefits to both patients and surgeons, including less postoperative pain and tissue damage, shorter hospital stay [hussain_use_2014, motkoski_toward_2013], reduced tremors compared to open surgeries [lanfranco_robotic_2004], and improved ergonomics compared to laparoscopic surgeries [kenngott_robotic_2008, lanfranco_robotic_2004].
@@
To obtain these advantages, robotic surgeons require substantial training [crawford2018evolution]. Although robotic surgical skill learning has been studied [narazaki2006robotic, hung2013comparative, ahmed2015development, yang2017effectiveness, abdelaal2018play, satava2020proving, chen2019objective, chang2003robotic, sinha2023current, rota2024implementation], there are currently many unknowns about how robotic surgeons acquire the necessary skills and a long-term study on surgical skill acquisition was needed. Therefore, in [sharon2025datasetanalysislongtermskill], surgical residents came three times throughout a shift once a month for six months. They performed several surgical training tasks; video and kinematic data were recorded to study their learning processes. All the data from was [sharon2025datasetanalysislongtermskill] released to enable further analyses. In the current work, we evaluate the performance in the ring tower transfer task, in which residents moved a ring along a curved wire as quickly as possible, while minimizing collisions with the wire.
@@
One challenge when assessing surgical skill is the method of evaluation. Currently, much of the evaluation is subjective, where one surgeon evaluates the performance of another and awards a score [vassiliou2005global, martin1997objective, goh2012global]. These scores suffer from large variability and require a great deal of time from the evaluating surgeons. RAMIS offers the possibility of overcoming these limitations by providing recorded kinematic and video data. This enables the automatic computation of metrics that are indicative of surgical skill [tausch2012content, kenney2009face, liang2018motion, sharon2021rate, chien2010accuracy, chen2019objective, boal2024evaluation, brentnall2024evaluation]. Metrics that have been used include task completion time [tausch2012content, liang2018motion, chien2010accuracy], path length [kenney2009face], and movement smoothness [liang2018motion]. In Sharon et al. [sharon2025datasetanalysislongtermskill], residents’ kinematic metrics, such as path length and rate of orientation change, [sharon2021rate] were computed for the ring tower transfer task to evaluate aspects of surgical skill.
@@
Accuracy, or error metrics, are additional important aspects of performance in many tasks [liang2018motion, chien2010accuracy], including surgery, where errors can result in patient injury [sarker2005errors, dankelman2005systems, bosma2011incidence, etchells2003patient]. The value of analyzing errors has been demonstrated in many motor control studies [fitts1954information, gallea2008error, reis2009noninvasive, shadmehr2010error]. Errors have been shown to facilitate different learning processes [herzfeld2014memory], distinguishing between implicit learning driven by error in the predicted outcome of movement [mazzoni2006implicit, taylor2011flexible, taylor2014explicit] and explicit learning driven by performance error [mcdougle2016taking, taylor2014explicit]. However, many of these motor control studies focus on simple tasks, and similar error metrics are challenging to compute in the more complex RAMIS tasks [chang2003robotic], leading many works to focus on other metrics, examples of which are listed above.
Developing algorithms for detecting surgical errors can serve as a step toward reducing surgical errors [shao2024think, sirajudeen2024deep, li2022runtime, xu2024sedmamba]. Furthermore, it can shed light on the learning processes of surgeons, eventually enabling the pinpointing of the different learning mechanisms that underlie skill acquisition in robotic surgical training. Moreover, detecting surgical errors can allow for designing error-related feedback for RAMIS, which has the potential to improve surgical performance [kluger1996effects, 11202637]. Strides have been made toward detecting surgical errors. In virtual simulations, such as Intuitive Surgical’s SimNow simulator, analysis of errors is now available [seeliger2024skill]. Furthermore, several works have developed algorithms for error detection in surgical tasks [shao2024think, sirajudeen2024deep, li2022runtime, xu2024sedmamba]. However, each of these studies has limitations, such as requiring a previously labeled dataset, and detecting errors in dry lab training remains an open challenge.
@@
In this work, we developed an image-processing algorithm for detecting collision errors in the ring tower transfer task. We used the video data from [sharon2025datasetanalysislongtermskill]  and wrote an algorithm to detect the collisions between the ring and the towers. Quantifying the errors can allow for studying their progression over the six-month experiment. As there is often a trade-off between speed and accuracy [chien2010accuracy, fitts1954information, reis2009noninvasive, listman2021long], we assessed completion time together with the errors to shed light on the surgical residents’ learning process of the task. Furthermore, this algorithm can be used to label the data for future analyses. The main contributions of our work are:
@@
    We discuss the residents’ learning process over six months based on completion time and error results.
@@
    We developed an algorithm to detect collision errors, enabling the error analyses. The general premise of the algorithm can potentially be adapted to detect collision errors in other tasks.
@@
    We will release our codes and labeled data so they can be used for future analyses.
@@ II-A Dataset
18 surgical residents with no robotic experience from the Soroka Medical Center, Israel visited the lab 18 times – before, during, and after a 26-hour shift approximately once a month for six months. In each visit, they used the da Vinci Research Kit (dVRK) [kazanzides2014open] to complete three surgical training tasks. They sat at the surgeon console and used the surgeon-side manipulators (SSMs) to teleoperate the patient-side manipulators (PSMs). The setup included two static Blackfly S cameras (FLIR Integrated Imaging Solutions Inc.) that captured the task board on the patient side. The video feed from the cameras was displayed on the stereo viewer in the surgeon console. The kinematic data, and video from the left camera, were recorded. In this work, we use the video data, which was presented to the residents at 35 Hz; however, the images were not recorded at a constant rate.
Here, we focus on the ring tower transfer task (Fig. 1); the residents extracted the black ring from the right vertical (RV) tower using their right hand, transferred it to their left hand, and inserted it onto the left horizontal (LH) tower. Next, they extracted the ring from the left vertical (LV) tower with their left hand and then inserted it onto the right horizontal (RH) tower using their right hand. The residents were asked to do so as fast as possible with minimal errors. Errors were defined as collisions between the tower and the ring. They received no feedback about their performance. Full details about the experiment and setup can be found in [sharon2025datasetanalysislongtermskill].
@@ II-B3 Error Detection
Each collision error of the ring with one of the towers was reflected as a movement of the tower in the recorded video. Therefore, to detect each error, our main steps were to: (1) locate the towers in the video frames, (2) compute all movement between each two consecutive frames, (3) sum the magnitude of the velocities of all pixels located as tower pixels, and (4) use a threshold to classify the error frames. Our goal was to detect these errors in order to analyze their progression over the 18 sessions. We will now elaborate on each step of the algorithm.
@@
Computation of all movement between each two consecutive frames: We used the MATLAB optical flow function with the Horn-Schunck method [horn1981determining]. From this, we obtained the magnitude of each of the image pixel velocities between every two consecutive images. This signal was likely affected by the lack of a constant sampling rate; however, it was still suitable for our needs.
@@
Lastly, we set several rules to clean the thresholded signal. The first was that a detected movement in the first 10 frames was only considered a movement if it continued also in the following five frames. This was because the first frame in every video was compared to a black image. Therefore, there was commonly a detected tower movement at the beginning, which was real in some cases, and an artifact in others. Next, all detected moves within 10 samples of each other were combined. This was to cope with cases in which the tower did not significantly move between two consecutive frames during an error. Third, all lone detected movement samples in a neighborhood of five samples were removed. Fourth, if there was a detected tower movement in the last 10 samples of an interaction with one of the vertical towers, the error lasted until the end of the interaction. We did not implement this for the horizontal towers, as the final part of the interaction was to place the ring on the tower, consistently leading to an intentional tower movement. Lastly, all cases of robot crashes were noted (this occurred a total of nine times during the interaction with one of the four towers), and the frames between the beginning of the crash and the return to the task were discarded from the data in all our analyses.
@@
Removal of End of Horizontal Tower Data: The residents placed the rings near the end of the horizontal towers. Hence, this was the only portion of the task in which a tower movement did not reflect an error, however, was still detected as one by the algorithm. Including this in the data introduces noise, and we therefore removed it as described in the following.
In an additional algorithm, we detected the black ring using color and size thresholds. Furthermore, as the ring is on the tower during the entire interaction with the tower, a minimal distance between each detected ring pixel and the nearest detected tower pixel was set. We do not elaborate on the details of this algorithm here, however, the codes and results are available in the GitHub repository.
To identify when the residents moved to the straight part of the horizontal tower, where the ring was placed, we set a limit on the horizontal axis. This was because collisions with the horizontal towers led to tower movements primarily in the vertical axis. We set several such limits, creating separate tower segments, the last of which was the straight part at the end where residents were asked to place the ring. As collisions in this segment do not reflect errors, rather the desired placement of the ring on the tower, we removed the data from this segment from the analyses.
Algorithm Analysis: Our algorithm outputted the frames with detected collisions for each resident’s interaction with each of the four towers. To check our algorithm, we displayed the results on videos. Our code used the automatically detected tower movements and created a copy of the recorded task video while coloring the tower red every time a tower movement was detected (Video 1). We found that our algorithm worked well for most of the collisions. To account for occasions of missed, or wrongly detected collision errors, we wrote a code enabling us to run through the video and correct the automatic labels with a simple button press when necessary. This manual check enabled us to quantify our algorithm’s ability to accurately detect the collision errors. To quantify its detection performance, we compared the frames detected as error frames by the algorithm with any corrections made. We computed the number of true positive (TP), true negative (TN), false positive (FP), and false negative (FN) frames for each tower and each resident.
Our algorithm achieved an accuracy of 94.79%, a TPR (true positive rate, the ability to detect an error when there is one) of 91.98%, a TNR (true negative rate, the ability to recognize that there is no error when there indeed is not a collision) of 98.93%, and an F1 score of 95.46%. These values demonstrate the high ability of the algorithm to accurately detect the collision errors. However, to allow for precise analysis of the residents’ learning processes, we used the corrected versions of the labels.
@@ II-B4 Surgical Learning Analysis
where $T$ is the time it took to complete the insertion or extraction from each tower in seconds. Each of these times was computed using the timestamps of the frames manually segmented as the start and end of the interaction with each tower. The timestamp of the start frame was subtracted from that of the end frame, and the times of the four towers were summed. Hence, the CompletionTimeCompletion\;Time reflects the interaction time with the four towers, however, unlike in [sharon2025datasetanalysislongtermskill] does not include the time spent transferring the rings between hands or the positioning at the start of each tower, as these data were not included in the current work.
@@
where $ET$ is the total time spent colliding with each tower. The error time for each tower was computed by subtracting the first timestamp of the error from the last timestamp of the error. These times were then summed for each tower, and then for all four towers. This total error time was then divided by the CompletionTimeCompletion\;Time to obtain the ErrorPercentageError\;Percentage.
@@ II-B5 Statistical Analyses
We used three two-way repeated measures ANOVA models to assess the progression of each of our three metrics between months and sessions. In each model, the independent variables were shift (categorical: 1-6), timing relative to the shift (categorical: before, during, after), and resident (random). We also tested the interaction between the shift and the timing relative to the shift to assess if the effect of the timing differed between the shifts. The dependent variables were each the CompletionTimeCompletion\;Time, NumberofErrorsNumber\;of\;Errors, and ErrorPercentageError\;Percentage in turn. For the CompletionTimeCompletion\;Time and NumberofErrorsNumber\;of\;Errors, we log-transformed the data to perform the analysis on normally distributed data. We used Mauchly’s test to check the assumption of sphericity and used Greenhouse-Geisser to adjust the degrees of freedom in the event of violations. The two participants with missing data were removed from the statistical analyses. We performed the statistical analyses using the MATLAB Statistics Toolbox. Statistical significance was determined at the 0.05 threshold.
@@ III Results
Using the detected errors, we computed the completion time and error metrics for all 18 residents (Fig. 3). Fig. 3(a) shows the individual results of all the residents, as well as the average trend across the 18 visits. We observed a significant decrease in CompletionTimeCompletion\;Time, both between shifts and the timing relative to the shifts (Table I). This is further shown in Fig. 3(b) and Fig. 3(c), which are averaged across the shift, and the timing relative to the shift, respectively. This decrease reflects that, on average, the surgical residents learned to complete the task faster as the experiment progressed. The interaction between the shift and timing relative to the shift was also significant, indicating that the improvement rates in the CompletionTimeCompletion\;Time between the three timings relative to the shift differed between the six shifts.
When examining the error metrics, we observed a relatively constant average ErrorPercentageError\;Percentage. There was no significant difference between shifts, reflecting that residents spent a similar percentage of time colliding with the towers in the six months. This demonstrates a lack of improvement in terms of reducing errors. There was a significant effect of the timing relative to the shift on the ErrorPercentageError\;Percentage, however, the effects were small (before: $65.12\pm 21.34\%$, during: $66.22\pm 22.56\%$, and after: $61.30\pm 23.03\%$). Lastly, the NumberofErrorsNumber\;of\;Errors significantly decreased between the six shifts, but not between the three timings relative to the shift. Taken together, the results show that the residents learned to complete the task faster, and began with making a larger number of shorter errors. With time, they made fewer separate errors, however, the portion of the time spent making the errors remained relatively constant. The statistical results are reported in Table I.
@@
Figure 3: Time and Error Results. (a) The results of all 18 residents in the 18 visits are shown in pale purple. The dark line shows the average, and the shaded area is the 95% confidence interval, of all the residents. (b) The average results of all the residents as a function of the shift number. The three shades of green show the three timings relative to the shift (before, during and after). (c) The average results of all the residents as a function of the timing relative to the shift. The six shades of blue show the six shifts. The top row shows the CompletionTimeCompletion\;Time, the middle row is the ErrorPercentageError\;Percentage and the bottom row is the NumberofErrorsNumber\;of\;Errors.
@@
Fig. 4 further illustrates the learning process in the time-error space. In terms of error metrics, we used the ErrorPercentageError\;Percentage, as it describes the amount of task time spent in error, whereas the NumberofErrorsNumber\;of\;Errors does not reflect their magnitude in the task. Fig. 4(a) shows the average ErrorPercentageError\;Percentage of all the residents as a function of the average CompletionTimeCompletion\;Time for each of the 18 visits. This illustrates the average improvement in the CompletionTimeCompletion\;Time with a lack of deterioration in the ErrorPercentageError\;Percentage.
@@
Fig. 4(b) shows each resident’s first and last ErrorPercentageError\;Percentage as a function of CompletionTimeCompletion\;Time, illustrating the individual process of each resident. We used a polynomial line to split the time-error space to demonstrate that all residents began on the right side of the line, and ended on the left. This line is similar to the Speed-Accuracy Tradeoff Functions shown in [reis2009noninvasive], characterizing the relation between the two. All the green symbols, representing the first visit, show a higher CompletionTimeCompletion\;Time relative to the blue symbols (last visit), illustrating that all of the residents learned to complete the task faster over the six months. Many of the residents also showed some improvement in the ErrorPercentageError\;Percentage, which taken together with the improvement in CompletionTimeCompletion\;Time shows overall learning and improvement in both the time and error aspects of the task. However, there were six residents who improved their CompletionTimeCompletion\;Time but increased their ErrorPercentageError\;Percentage.
@@
Figure 4: Time-Error Analysis. (a) The average ErrorPercentageError\;Percentage as a function of the average CompletionTimeCompletion\;Time across all 18 residents. The green shades indicate earlier sessions, and blue shades are the later ones, as shown in the color bar. (b) The ErrorPercentageError\;Percentage as a function of CompletionTimeCompletion\;Time for each of the 18 residents in the first (green) and last (blue) visit. Each resident is shown by a different marker, and the first and last visits of each resident are connected with a line. To further demonstrate the improvement, the gray polynomial line splits the time-error space, such that all residents began on the right of the line and ended on the left.